# Cosmological constraints from the Minkowski functionals of the BOSS CMASS galaxy sample

### v1

@@ 3.6.1 Priors
We use DYNESTY [147] to estimate Bayesian posteriors using the dynamic nested sampling methods [148], with the likelihood defined as Eq 3.9, the prior function given in Eq 3.19, a Gaussian prior on $\omega_{\rm b}$ given in Eq 3.18, and flat priors in the range listed in table 1 on other parameters. For our baseline likelihood analysis, the four extensions of the cosmological parameters $\{\alpha_{s},N_{\rm eff},w_{0},w_{a}\}$ are fixed to $\{0.0,3.0146,-1.0,0.0\}$, corresponding to the $\Lambda$CDM limit. Most of the results presented in this work are obtained with this default pipeline, but we also explore other options in this work.
@@ 5.1 Fits of the summary statistics
In figure 4, we compare the emulator predictionsuwith the CMASS measurement. Overall, we find the emulator agrees best with the data for the MFs (2PCF) when using the best-fit value from the MFs (2PCF). However, the agreement is weakened when feeding the best-fit value from the MFs (2PCF) to the 2PCF (MFs) emulator. This is understandable since the MFs and 2PCF probe different properties of the LSS. In addition, the MFs focus on structures with scales around the smoothing scale while the 2PCF measures the clustering of galaxies for a wide range of separations. Interestingly, we find large deviations of the MFs emulator from the CMASS data for large thresholds when inputting the best-fit value from the 2PCF. For these high $\delta$ values, the non-Gaussianity becomes important and the 2PCF fails to capture the non-Gaussian information. On the other hand, the Gaussian smoothing used for MFs smears out very small-scale information. That is why the 2PCF emulator can’t accurately reproduce both $\xi_{0}$ and $\xi_{2}$ of the CMASS data for $s\lesssim 10h^{-1}\rm Mpc$ when inputting the best-fit value from the MFs.
@@
We report the goodness-of-fit parameters in table 2 for the best fits to the CMASS data. When the covariance matrix only includes the contribution from the CMASS sample variance ($C_{\rm data}$), we obtained values 1.05, 1.25, and 1.63 for the 2PCF, MFs, and their combination. The value for the 2PCF is comparable to those reported in [112, 59, 75], where $\chi^{2}/dof$ was found to be in the range of 1.06 to 1.15. The value for the MFs is smaller than that for the density split clustering ($\chi^{2}/dof=1.67$ [59]) and also for the wavelet scattering transform ($\chi^{2}/dof=1.37$ [75]). However, the goodness-of-fit values for the summary statistics and their combination are all smaller than one once including the error budget from the emulator error ($C_{\rm data}+C_{\rm emu}$). This indicates an overestimate of the emulator error, which may be more properly estimated with the Abacus mocks whose summary statistics are close to those of the CMASS data. A similar approach was adopted in [112], where the error is estimated using only HOD catalogs whose likelihood is high with respect to the CMASS measurement. Our estimate of $C_{\rm emu}$ is more conservative because we have included the error around regions near the edge of the priors, where the emulator is most likely to fail to correctly study the statistics. The goodness-of-fit values are further reduced after the addition of the sample variance for the Abacus mocks ($C_{\rm data}+C_{\rm emu}+C_{\rm abacus}$), which is a suggestion that the double counting of sample variance in the CMASS data and the Abacus mocks may lead to an overestimate of error in the analysis. We will investigate this issue in future work, and we note the phase correction routine used by [112, 75] may provide a method to alleviate this problem.
@@ 5.2 Constraints on $\Lambda$CDM
Compared with the 2PCF, the MFs provide a factor of 2.0, 1.7, 1.4, and 2.0 times tighter constraints on $\omega_{\rm cdm}$, $\sigma_{8}$, $n_{s}$, and $h$, respectively. The amplitudes of the MFs are sensitive to the shape of the linear matter power spectrum and thus to $\omega_{\rm cdm}$ and $n_{s}$ [83, 34, 35]. In redshift space, the amplitudes are also sensitive to the growth rate of the structure (thus $\omega_{\rm cdm}$) and galaxy bias (HOD parameters)[90, 91]. They are practically insensitive to the amplitude of the power spectrum, that is, $\sigma_{8}$. However, the shapes of the MFs are very sensitive to $\sigma_{8}$ because it controls the amplitude of the departure from the Gaussian shape of the MFs and determines the position and scaling of the shape through the threshold parameter $\delta$ [91, 81]. Extra constraining power on $\omega_{\rm cdm}$ and $n_{s}$ comes from the non-Gaussian information extracted by the MFs, as found in [78, 81]. Otherwise, the MFs are anticipated to provide similar constraints (theoretically) on the two parameters to those from the 2PCF because there is only Gaussian information from the amplitudes of the MFs. The strong constraint on $h$ may stem from the $\theta_{*}$ prior because we do not expect to capture more accurate information of $h$ than the 2PCF with a single smoothing scale, which is similar to what was found in [75].
@@ Acknowledgments
We would like to express our heartfelt gratitude to Jeremy L. Tinker for generously generating the Uchuu mock for us. We are grateful to Zhongxu Zhai for kindly helping us transfer the Uchuu mock data. WL also thanks Zhongxu Zhai and Cheng Zhao for helpful discussions. This work is supported by the National Natural Science Foundation of China Grants No. 12173036 and 11773024, by the National Key R&D Program of China Grant No. 2021YFC2203100 and No. 2022YFF0503404, by the China Manned Space Project “Probing dark energy, modified gravity and cosmic structure formation by CSST multi-cosmological measurements” and Grant No. CMS-CSST-2021-B01, by the Fundamental Research Funds for Central Universities Grants No. WK3440000004 and WK3440000005, by Cyrus Chun Ying Tang Foundations, and by the 111 Project for "Observational and Theoretical Research on Dark Matter and Dark Energy" (B23042). EP was in part supported by the U.S. Department of Energy HEP-AI program grant DE-SC0023892.
@@ Appendix C Constraints of HOD parameters
We plot the posterior distribution of the full cosmological and HOD parameters from the 2PCF, the MFs, and the combination of the 2PCF and MFs for the BOSS CMASS galaxy sample in figure 15. The corresponding best-fit, mean, and $68\%$ confidence interval values of the HOD parameters are also listed in table 5. We find the MFs are more sensitive to all of the HOD parameters except for the velocity bias parameters $\alpha_{vel,c}$ and $\alpha_{vel,s}$. There are two reasons for this. On one hand, a considerable proportion of the velocity information related to the velocity bias parameters is smeared out by the isotropic Gaussian smoothing process since we have used a relatively large smoothing scale $R_{G}=15~{}h^{-1}\rm Mpc$. Therefore, if the change of velocity bias parameter only leads to variations in the LSS with scales much smaller than the smoothing scale, the MFs can fail to detect its impact and the sensitivity of the MFs to this change is low. We find the constraining power of the MFs on the two velocity bias parameters can be increased when small smoothing scales are used. On the other hand, the MFs collect contributions from the infinitesimal volume or area element in an angle-independent way, as seen from equation 3.4, this means the anisotropic information can be buried when adding up contributions that involve the LoS direction and those that involve directions orthogonal to the LoS [91, 80]. Therefore, when the lost anisotropic information is added by the 2PCF (in particular, $\xi_{2}$), we find the constraints from the MFs on $\alpha_{vel,c}$ and $\alpha_{vel,c}$ are improved by a factor of 2.3 and 2.4, respectively.
@@
We find our constraints on $logM_{cut}$, $logM_{1}$, $log\sigma$, and $\alpha$ are in agreement with the baseline results of [114]. However, their measurement for $\kappa$ is much lower than our inference. The joint analysis of the two statistics infers $\alpha_{vel,c}=0.19\pm 0.10$ and $\alpha_{vel,s}=1.02\pm 0.16$, that is, a nonzero central velocity bias with $\sim 2\sigma$ significance and a satellite velocity bias consistent with 1. This measurement is consistent with the results of [114, 118], both of the two works found $\alpha_{vel,c}\simeq 0.2$ and $\alpha_{vel,c}\simeq 1.0$. We note that the galaxy assembly bias is not implemented in the analysis, which will be investigated in our future work.

### v2

@@ 3.6.1 Priors
We use DYNESTY [146] to estimate Bayesian posteriors using the dynamic nested sampling methods [147], with the likelihood defined as Eq 3.9, the prior function given in Eq 3.19, a Gaussian prior on $\omega_{\rm b}$ given in Eq 3.18, and flat priors in the range listed in table 1 on other parameters. For our baseline likelihood analysis, the four extensions of the cosmological parameters $\{\alpha_{s},N_{\rm eff},w_{0},w_{a}\}$ are fixed to $\{0.0,3.0146,-1.0,0.0\}$, corresponding to the $\Lambda$CDM limit. Most of the results presented in this work are obtained with this default pipeline, but we also explore other options in this work. The priors used in this work are summarised in Table 2.
Priors Descriptions
Gaussian prior for $\omega_{\mathrm{b}}$ $\omega_{\mathrm{b}}=0.02268\pm 0.00038$ ($\omega_{\mathrm{b}}=0.02230\pm 0.00038$ for test on Uchuu)
Flat prior except $\omega_{\mathrm{b}}$ Uniform prior distribution with the same bounds listed in Table 1
Number density constraint $p(\theta)=(H[\bar{n}(\theta)-0.00024]-1)\times\infty$
Table 2: Summary of priors used in this work.
@@ 5.1 Fits of the summary statistics
In figure 4, we compare the emulator predictions with the CMASS measurement. Overall, we find the emulator agrees best with the data for the MFs (2PCF) when using the best-fit value from the MFs (2PCF). However, the agreement is weakened when feeding the best-fit value from the MFs (2PCF) to the 2PCF (MFs) emulator. This is understandable since the MFs and 2PCF probe different properties of the LSS. In addition, the MFs focus on structures with scales around the smoothing scale while the 2PCF measures the clustering of galaxies for a wide range of separations. Interestingly, we find large deviations of the MFs emulator from the CMASS data for large thresholds when inputting the best-fit value from the 2PCF. For these high $\delta$ values, the non-Gaussianity becomes important and the 2PCF fails to capture the non-Gaussian information. On the other hand, the Gaussian smoothing used for MFs smears out very small-scale information. That is why the 2PCF emulator can’t accurately reproduce both $\xi_{0}$ and $\xi_{2}$ of the CMASS data for $s\lesssim 10h^{-1}\rm Mpc$ when inputting the best-fit value from the MFs.
@@
We report the goodness-of-fit parameters in table 3 for the best fits to the CMASS data. When the covariance matrix only includes the contribution from the CMASS sample variance ($C_{\rm data}$), we obtained values 1.53, 1.25, and 1.63 for the 2PCF, MFs, and their combination. The value for the 2PCF is larger than those reported in [112, 59, 75], where $\chi^{2}/dof$ was found to be in the range of 1.06 to 1.15. Nevertheless, we believe our value is reasonable considering the significant tension between the BOSS 2PCF monopole and our best fit on large scales seen in figure 4. The value for the MFs is smaller than that for the density split clustering ($\chi^{2}/dof=1.67$ [59]) and also for the wavelet scattering transform ($\chi^{2}/dof=1.37$ [75]). However, the goodness-of-fit values for the summary statistics and their combination are all smaller than one once including the error budget from the emulator error ($C_{\rm data}+C_{\rm emu}$). This indicates an overestimate of the emulator error, which may be more properly estimated with the Abacus mocks whose summary statistics are close to those of the CMASS data. A similar approach was adopted in [112], where the error is estimated using only HOD catalogs whose likelihood is high with respect to the CMASS measurement. Our estimate of $C_{\rm emu}$ is more conservative because we have included the error around regions near the edge of the priors, where the emulator is most likely to fail to correctly study the statistics. The goodness-of-fit values are further reduced after the addition of the sample variance for the Abacus mocks ($C_{\rm data}+C_{\rm emu}+C_{\rm abacus}$), which is a suggestion that the double counting of sample variance in the CMASS data and the Abacus mocks may lead to an overestimate of error in the analysis. We will investigate this issue in future work, and we note the phase correction routine used by [112, 75] may provide a method to alleviate this problem.
@@ 5.2 Constraints on $\Lambda$CDM
Compared with the 2PCF, the MFs provide a factor of 2.0, 1.7, 1.4, and 2.0 times tighter constraints on $\omega_{\rm cdm}$, $\sigma_{8}$, $n_{s}$, and $h$, respectively. This is consistent with [168], where they found the MFs of weak lensing convergence maps also have stronger constraining power on $\Omega_{m}$ and $S_{8}=\sigma_{8}\sqrt{\Omega_{m}/0.3}$ than two-point statistics. The amplitudes of the MFs are sensitive to the shape of the linear matter power spectrum and thus to $\omega_{\rm cdm}$ and $n_{s}$ [83, 34, 35]. In redshift space, the amplitudes are also sensitive to the growth rate of the structure (thus $\omega_{\rm cdm}$) and galaxy bias (HOD parameters) [90, 91]. They are practically insensitive to the amplitude of the power spectrum, that is, $\sigma_{8}$. However, the shapes of the MFs are very sensitive to $\sigma_{8}$ because it controls the amplitude of the departure from the Gaussian shape of the MFs and determines the position and scaling of the shape through the threshold parameter $\delta$ [91, 81]. Extra constraining power on $\omega_{\rm cdm}$ and $n_{s}$ comes from the non-Gaussian information extracted by the MFs [^19] The non-Gaussian information is not necessarily extracted only from the departure from the Gaussian shape of the MFs, because the amplitude of the MFs also receives contributions from the trispectrum [83, 91, 84]., as found in [78, 81]. Otherwise, the MFs are anticipated to provide similar constraints (theoretically) on the two parameters to those from the 2PCF. The strong constraint on $h$ may stem from the $\theta_{*}$ prior because we do not expect to capture more accurate information of $h$ than the 2PCF with a single smoothing scale, which is similar to what was found in [75].
@@ Acknowledgments
We would like to express our heartfelt gratitude to Jeremy L. Tinker for generously generating the Uchuu mock for us. We are grateful to Zhongxu Zhai for kindly helping us transfer the Uchuu mock data. WL also thanks Zhongxu Zhai, Cheng Zhao, and Aoxiang Jiang for helpful discussions. We are also grateful to the anonymous referee for the constructive comments, which helped to improve the quality of this paper. This work is supported by the National Natural Science Foundation of China Grants No. 12173036 and 11773024, by the National Key R&D Program of China Grant No. 2021YFC2203100 and No. 2022YFF0503404, by the China Manned Space Program with Grant No. CMS-CSST-2025-A04 and Grant No. CMS-CSST-2021-B01, by the Fundamental Research Funds for Central Universities Grants No. WK3440000004 and WK3440000005, by Cyrus Chun Ying Tang Foundations, and by the 111 Project for "Observational and Theoretical Research on Dark Matter and Dark Energy" (B23042). EP was in part supported by the U.S. Department of Energy HEP-AI program grant DE-SC0023892.
@@ Appendix C Constraints of HOD parameters
We plot the posterior distribution of the full cosmological and HOD parameters from the 2PCF, the MFs, and the combination of the 2PCF and MFs for the BOSS CMASS galaxy sample in figure 15. The corresponding best-fit, mean, and $68\%$ confidence interval values of the HOD parameters are also listed in table 6. We find the MFs are more sensitive to all of the HOD parameters except for the velocity bias parameters $\alpha_{vel,c}$ and $\alpha_{vel,s}$. There are two reasons for this. On one hand, a considerable proportion of the velocity information related to the velocity bias parameters is smeared out by the isotropic Gaussian smoothing process since we have used a relatively large smoothing scale $R_{G}=15~{}h^{-1}\rm Mpc$. Therefore, if the change of velocity bias parameter only leads to variations in the LSS with scales much smaller than the smoothing scale, the MFs can fail to detect its impact and the sensitivity of the MFs to this change is low. We find the constraining power of the MFs on the two velocity bias parameters can be increased when small smoothing scales are used. On the other hand, the MFs collect contributions from the infinitesimal volume or area element in an angle-independent way, as seen from equation 3.4, this means the anisotropic information can be buried when adding up contributions that involve the LoS direction and those that involve directions orthogonal to the LoS [91, 80]. Therefore, when the lost anisotropic information is added by the 2PCF (in particular, $\xi_{2}$), we find the constraints from the MFs on $\alpha_{vel,c}$ and $\alpha_{vel,c}$ are improved.
@@
We find our constraints on $logM_{cut}$, $logM_{1}$, $log\sigma$, and $\alpha$ are in agreement with the baseline results of [114]. However, their measurement for $\kappa$ is much lower than our inference. The joint analysis of the two statistics infers $\alpha_{vel,c}=0.19\pm 0.10$ and $\alpha_{vel,s}=1.02\pm 0.16$, that is, a nonzero central velocity bias with $\sim 2\sigma$ significance and a satellite velocity bias consistent with 1. This measurement is consistent with the results of [114, 118], both of the two works found $\alpha_{vel,c}\simeq 0.2$ and $\alpha_{vel,c}\simeq 1.0$. We note that the galaxy assembly bias is not implemented in the analysis, which will be investigated in our future work. The mass resolution of the fiducial cosmology in the AbacusSummit simulation [105] is $2\times 10^{9}h^{-1}\mathrm{M}{\odot}$. Since the mass resolution is proportional to $\Omega_{m}$, and the varying range of $\Omega_{m}$ is approximately ±$\pm 20\%$, the simulations for non-fiducial cosmologies maintain a similar mass resolution. The best-fit values for $logM_{cut}$ and $logM_{1}$, obtained from the combination of the 2PCF and MFs, are 12.77 and 13.79, respectively. Therefore, the mass resolution of the AbacusSummit simulation is more than sufficient to resolve the lightest halos hosting central and satellite galaxies for the CMASS galaxy sample.
To study the effect of $\alpha_{cel,s}$ and $\alpha_{cel,c}$ on the MFs and disentangle their effects from other HOD parameters such as $logM_{cut}$ and $logM_{1}$, we plot in figure 16 the MFs of the HOD mocks in both real and redshift space, with and without central or satellite velocity bias, and for different values of $logM_{cut}$ or $logM_{1}$. We compare the difference in the MFs between real space and redshift space here because the effect of velocity bias is related to redshift space distortions (see [80] for a detailed interpretation of the effect of RSD on the MFs). We find the effect of RSD on the MFs is reduced when $\alpha_{vel,c}=0.8$  [^21] This phenomenon becomes more pronounced when we increase $\alpha_{vel,c}$ further. However, in this case, we aim to show the differences in the MFs caused by parameter variations within the hard prior range used in this analysis. or $\alpha_{vel,s}=1.0$. In these cases, the relative motions of both central and satellite galaxies with respect to their host halos cause the density field in redshift space to resemble that in real space to some extent. This finding is consistent with [118], where Fig. 6 shows that a low satellite velocity bias tends to make the contours more squashed along the line-of-sight direction, while a high central velocity bias has the opposite effect. Additionally, we observe a (weak) anticorrelation between $\alpha_{vel,c}$ and $\alpha_{vel,s}$ in the bottom right corner of our Figure 15, similar to what is seen in their Fig. 5. This anticorrelation arises because $\alpha_{vel,c}$ and $\alpha_{vel,s}$ can partially compensate for each other.
On the other hand, $logM_{cut}$ and $logM_{1}$ affect the MFs in a different manner. The most notable feature is that the MFs curves either expand or compress as the values of these two parameters change. Since $logM_{cut}$ controls the halo mass required to host a central galaxy, a larger $logM_{cut}$ corresponds to a higher linear galaxy bias. As a result, we observe more extended MFs curves for $logM_{cut}=13.3$ compared to $logM_{cut}=12.4$. The effect of $logM_{1}$ is more intricate: although smaller $logM_{1}$ allows halos with lower masses to host satellite galaxies, larger halos will also accommodate more satellite galaxies (since the power-law index for the number of satellite galaxies, $\alpha$, is larger than one for the CMASS galaxies). This means that a larger fraction of galaxies can be hosted by heavier halos in the HOD mock with smaller $logM_{1}$. Consequently, it is understandable that the MFs curves exhibit a higher galaxy bias for $logM_{1}=13.0$ than for $logM_{1}=15.0$. $logM_{cut}$ and $logM_{1}$ can also affect the amplitude of the MFs, as they influence the shape of the power spectrum (see Section 2 of [177] for an analytical description of how these parameters impact the power spectrum).
We hope that the explanation of how the four HOD parameters influence the MFs in different ways will help readers understand how the MFs can infer the velocity bias parameters, along with the other HOD parameters that determine galaxy bias.
Figure 16: The MFs of the HOD mock in real and redshift space (first column), with and without central (second column) or satellite velocity bias (third column), with different values of $logM_{cut}$ (fourth column) or $logM_{1}$ (last column). The values of other HOD parameters are set to the best value from the MFs.